The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall Inc. common stock was $ 62 on December 31, 20Y2. Marshall Inc. Comparative Retained Earnings Statement For the Years Ended December 31, 20Y2 and 20Y1 20Y2 20Y1 Retained earnings, January 1 $ 1,652,575 $ 1,392,225 Net income 379,600 285,100 Total $ 2,032,175 $ 1,677,325 Dividends On preferred stock $ 6,300 $ 6,300 On common stock 18,450 18,450 Total dividends $ 24,750 $ 24,750 Retained earnings, December 31 $ 2,007,425 $ 1,652,575 Marshall Inc. Comparative Income Statement For the Years Ended December 31, 20Y2 and 20Y1 20Y2 20Y1 Sales $ 2,433,455 $ 2,242,090 Cost of goods sold 854,100 785,770 Gross profit $ 1,579,355 $ 1,456,320 Selling expenses $ 553,570 $ 675,860 Administrative expenses 471,555 396,940 Total operating expenses 1,025,125 1,072,800 Income from operations $ 554,230 $ 383,520 Other income 29,170 24,480 $ 583,400 $ 408,000 Other expense (interest) 152,000 84,000 Income before income tax $ 431,400 $ 324,000 Income tax expense 51,800 38,900 Net income $ 379,600 $ 285,100 Marshall Inc. Comparative Balance Sheet December 31, 20Y2 and 20Y1 Dec. 31, 20Y2 Dec. 31, 20Y1 Assets Current assets Cash $ 390,150 $ 353,670 Marketable securities 590,490 586,080 Accounts receivable (net) 423,400 401,500 Inventories 321,200 248,200 Prepaid expenses 73,809 70,730 Total current assets $ 1,799,049 $ 1,660,180 Long-term investments 1,388,059 508,028 Property, plant, and equipment (net) 2,090,000 1,881,000 Total assets $ 5,277,108 $ 4,049,208 Liabilities Current liabilities $ 599,683 $ 576,633 Long-term liabilities Mortgage note payable, 8 % $ 850,000 $ 0 Bonds payable, 8 % 1,050,000 1,050,000 Total long-term liabilities $ 1,900,000 $ 1,050,000 Total liabilities $ 2,499,683 $ 1,626,633 Stockholders' Equity Preferred $ 0.70 stock, $ 40 par $ 360,000 $ 360,000 Common stock, $ 10 par 410,000 410,000 Retained earnings 2,007,425 1,652,575 Total stockholders' equity $ 2,777,425 $ 2,422,575 Total liabilities and stockholders' equity $ 5,277,108 $ 4,049,208 Required: Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirement, if required. Assume 365 days a year. 1. Working capital $ 1,199,366 2. Current ratio 3 3. Quick ratio 2.3 4. Accounts receivable turnover 5.9 5. Number of days' sales in receivables 61.9 days 6. Inventory turnover 3 7. Number of days' sales in inventory 121.7 days 8. Ratio of fixed assets to long-term liabilities 1.1 9. Ratio of liabilities to stockholders' equity 0.9 10. Times interest earned 3.8 11. Asset turnover 12. Return on total assets % 13. Return on stockholders’ equity % 14. Return on common stockholders’ equity % 15. Earnings per share on common stock $ 16. Price-earnings ratio 17. Dividends per share of common stock $ 18. Dividend yield %
Answer #1

Hi,

It is humble request to provide data in appropriate format which helps in quick response. Here data is provided all in texts and not in tables and hence at least an hour have gone into generating basic data as under :

Marshall Inc.
Comparative Income Statement For the Years Ended December 31, 20Y2 December 31, 20Y1
Sales 2,433,455 2,242,090
Cost of goods sold 854,100 785,770
Gross profit 1,579,355 1,456,320
Selling expenses 553,570 675,860
Administrative expenses 471,555 396,940
Total operating expenses 1,025,125 1,072,800
Income from operations 554,230 383,520
Other income 29,170 24,480
Earning before Interest & Tax 583,400 408,000
Other expense (interest) 152,000 84,000
Income before income tax 431,400 324,000
Income tax expense 51,800 38,900
Net income 379,600 285,100
Marshall Inc.
Comparative Balance Sheet December 31, 20Y2 December 31, 20Y1
Assets
Current assets
Cash 390,150 353,670
Marketable securities 590,490 586,080
Accounts receivable 423,400 401,500
Inventories 321,200 248,200
Prepaid expenses 73,809 70,730
Total current assets 1,799,049 1,660,180
Long-term investments 1,388,059 508,028
Property, plant, and equipment (net) 2,090,000 1,881,000
Total assets 5,277,108 4,049,208
Liabilities
Current liabilities 599,683 576,633
Long-term liabilities
Mortgage note payable, 8 % 850,000 0
Bonds payable, 8 % 1,050,000 1,050,000
Total long-term liabilities 1,900,000 1,050,000
Total liabilities 2,499,683 1,626,633
Stockholders' Equity
Preferred $ 0.70 stock, $ 40 par 360,000 360,000
Common stock, $ 10 par 410,000 410,000
Retained earnings 2,007,425 1,652,575
Total stockholders' equity 2,777,425 2,422,575
Total liabilities and stockholders' equity 5,277,108 4,049,208
Marshall Inc.
Comparative Retained Earnings Statement For the Years Ended December 31, 20Y2 December 31, 20Y1
Retained earnings 1,652,575 1,392,225
Net income 379,600 285,100
Dividends On preferred stock -6,300 -6,300
Dividned On common stock -18,450 -18,450
Retained earnings 2,007,425 1,652,575

Now few ratios were given and few were required to be calculated. I have provided formula and numerical working for all the ratio including those for which answer was already available :

Working capital 1,199,366 Current Assets - Current Liabilities =1,799,049-599,683
Current ratio 3.00 Current Assets/ Current Liabilities=1,799,049/599,683
Quick ratio 2.34 (Current Asset-Inventory-Prepaid Expenses) / Curren+C43t Liabilities= (1,799,049-321,200-73809)/599,683
Accounts receivable turnover 5.9 Sales/(Opening Account Receivable+Closing Account Receivable)/2) =2,433,455/((401,500+423,400)/2)
Number of days' sales in receivables 61.9 (Opening Account Receivable+Closing Account Receivable)/2)/sale*365 = ((401,500+423,400)/2)2,433,455*365
Inventory turnover 3 Cost of Goods Sold/ ((Opening Inventory+Closing Inventory)/2) = 854,100/((321,200+248,200)/2)
Number of days' sales in inventory (Days) 121.7 Days in year / Inventory Turnover Ratio = 365/3
Ratio of fixed assets to long-term liabilities 1.1 Fixed Asset / Long Term Liabilites = 2,090,000/1,900,000
Ratio of liabilities to stockholders' equity 0.9 Total Liabilities / Equity = 2,499,683/2,777,425
Times interest earned 3.8 Earning before interest & Tax / Interest = 583,400/152,000
Asset turnover 0.5 Sales/(Beginning total asset+closing total asset)/2)=2,433,455/((4,049,208+5,277,108)/2)
Return on total assets % 0.13 Earning before interest & Tax / (Beginning total asset+closing total asset)/2)=583,400/((4,049,208+5,277,108)/2)
Return on stockholders’ equity % 0.49 Earning after Tax /(Common stock+Preferred stock) = 379,700/(410,000+360,000)
Return on common stockholders’ equity % 0.91 (Earning after tax-Dividend on preferred stock)/Common Stock = (379,700-6,300)/410,000
Earnings per share on common stock 9.10 (Earning after tax-Dividend on preferred stock)/Common Stock number = (379,700-6,300)/410,000/10
Price-earnings ratio 6.81 Market Price/Earning per share on common Stock = 62/9.10
Dividends per share of common stock 0.45 Dividend on Common stock / Number of Common stock = 18450/41000

Dividend yield %

0.05 Dividend per share on common stock /Price per common stock = 0.45/10

Hope this is helpful - do provide positive feedback.